Deacon Company is a merchandising company that is preparing a budget for the three-month period ended June 30th. The following information is available

Deacon Company
Balance Sheet
March 31
Assets
Cash $ 55,000
Accounts receivable 36,000
Inventory 40,000
Buildings and equipment, net of depreciation 100,000
Total assets $ 231,000
Liabilities and Stockholders’ Equity
Accounts payable $ 51,300
Common stock 70,000
Retained earnings 109,700
Total liabilities and stockholders’ equity $ 231,000
Budgeted Income Statements
April May June
Sales $ 100,000 $ 110,000 $ 130,000
Cost of goods sold 60,000 66,000 78,000
Gross margin 40,000 44,000 52,000
Selling and administrative expenses 15,000 16,500 19,500
Net operating income $ 25,000 $ 27,500 $ 32,500

Budgeting Assumptions:

  1. 60% of sales are cash sales and 40% of sales are credit sales. Twenty percent of all credit sales are collected in the month of sale and the remaining 80% are collected in the month subsequent to the sale.

  2. Budgeted sales for July are $140,000.

  3. 10% of merchandise inventory purchases are paid in cash at the time of the purchase. The remaining 90% of purchases are credit purchases. All purchases on credit are paid in the month subsequent to the purchase. The accounts payable at March 31 will be paid in April.

  4. Each month’s ending merchandise inventory should equal $10,000 plus 50% of the next month’s cost of goods sold.

  5. Depreciation expense is $1,000 per month. All other selling and administrative expenses are paid in full in the month the expense is incurred.

Required:

Prepare a budgeted balance sheet at June 30th. (Hint: You need to calculate the cash paid for selling and administrative expenses during April, May, and June to determine the cash balance in your June 30th balance sheet.)

Deacon Company
Balance Sheet
June 30
Assets
Cash ?
Accounts receivable 41,600
Inventory 52,000
Buildings and equipment, net of depreciation 97,000
Total assets $190,600
Liabilities and Stockholders’ Equity
Accounts payable $72,900
Retained earnings ?
Total liabilities and stockholders’ equity $72,900
Answer #1

Calculation of cash balance at end :

cash collection from receivables = Beginning receivables +total sales during the quarter-ending receivables

                     = 36000 + 340000 -41600

                      = 334400

Inventory Purchase = Total of cost of goods sold during quarter+ending inventory -beginning inventory

                     = 204000 +52000-40000

                      = 216000

Cash payment for inventory purchase =Beginning accounts payable+ Inventory purchase- ending accounts payable

                   = 51300 +216000 -72900

                    = 194400

cash payment for selling and administrative expense =Total selling and administrative expense -total depreciation

                           = 51000 -3000

                          = 48000

Cash Balance at end =Beginning balance +cash receipts -cash payments

             = 55000 +334400 - 194400 -48000

            = 147000

Total asset = 147000+41600+52000+97000= 337600

2)

Total Net income =25000+27500+32500=85000

Retained earning :Beginning balance +net income

             = 109700 +85000

                 = 194700

Total liabilities and stockholders equity = 72900+194700+70000

                               = 337600